OpenStudy (anonymous):

solve by completing the square. x^2+2x-7=0

OpenStudy (anonymous):

\[(b-\sqrt{b ^{2}-4ac})/2a\]

OpenStudy (anonymous):

the coefficients infront of every compomnet of the equation is the a,b,and c

OpenStudy (solomonzelman):

x^2+2x-7=0 add 13 to both sides x^2+2x-7+13=0+13 x^2+2x+4=13 (x+2)^2=13